# Defending Against Unforeseen Failure Modes with Latent Adversarial Training

### v3

@@ Abstract
Despite extensive diagnostics and debugging by developers, AI systems sometimes exhibit harmful unintended behaviors. Finding and fixing these is challenging because the attack surface is so large – it is not tractable to exhaustively search for inputs that may elicit harmful behaviors. Red-teaming and adversarial training (AT) are commonly used to improve robustness, however, they empirically struggle to fix failure modes that differ from the attacks used during training. In this work, we utilize latent adversarial training (LAT) to defend against vulnerabilities without generating inputs that elicit them. LAT leverages the compressed, abstract, and structured latent representations of concepts that the network actually uses for prediction. We use it to remove trojans and defend against held-out classes of adversarial attacks. We show in image classification, text classification, and text generation tasks that LAT usually improves both robustness to novel attacks and performance on clean data relative to AT. This suggests that LAT can be a promising tool for defending against failure modes that are not explicitly identified by developers. [^1] Code is available at https://github.com/thestephencasper/latent_adversarial_training.
@@ 1 Introduction
Ensuring that AI systems will be trustworthy, even in the face of anomalous and adversarial inputs, has been a major focus of research in the past decade (Szegedy et al., 2013; Goodfellow et al., 2014; Zhao et al., 2022), and it has been incorporated into risk management frameworks for AI governance (NIST, 2023; DSIT, 2023; EU, 2021; NISSTC, 2023). Developers commonly use test sets, red-teaming, and attack methods to identify vulnerabilities followed by adversarial training (AT) to fix them. This is valuable, but sometimes fails to address problems. There are often systematic differences between the failure modes that developers search for (e.g., $L_{p}$-norm attacks) and ones that models can exhibit post-deployment (e.g. (Hendrycks et al., 2021a, c)). Many real-world vulnerabilities can evade detection such as trojans (Hubinger et al., 2024; Carlini et al., 2022), jailbreaks (Liu et al., 2023; Wei et al., 2023; Zou et al., 2023b; Shah et al., 2023) novel attacks (Brown et al., 2018; Shayegani et al., 2023; Geiping et al., 2024), or black swans (Kolt, 2023; Hendrycks et al., 2021b). In Figure 1a, we illustrate the gap between failure modes developers identify, and the “unforeseen” ones they do not.
@@
Our motivation for LAT is based on the vastness of a model’s input space compared to the manifold of task-relevant features within it. Across the latents, a model gradually develops more compressed, abstract, and structured representations of the concepts it uses to process information. [^2] Since the mappings from layer to layer are not typically surjective, not all latent states are inducible by real inputs. This relaxation of the problem is what gives LAT the potential to compensate for the gap between failure modes that are identified by developers and ones that are not (see Figure 1a). This makes it possible for latent space attacks to activate neural circuitry that elicits failures (Figure 1b) without requiring inputs that trigger them. Thus, we hypothesize that even if it is difficult to find a problem with a model using input space attacks, it may be comparatively easier with latent space ones. We make three contributions:
Figure 1: (a) We study latent adversarial training (LAT) as a method to reduce risks from failure modes that are not found by developers pre-deployment. (b) Our motivation for LAT is based on how models develop more compressed, abstract, and structured representations across their latents. We hypothesize that many failures that are difficult to elicit from the input space may be easier to elicit from the latent space. (c) In experiments, we compare AT with LAT. The key difference is that AT applies adversarial perturbations to the input, and LAT applies adversarial perturbations to a hidden layer. (d) In each of our experiments, we compare methods based on (1) their performance on clean evaluation data, (2) their robustness to novel classes of adversarial examples not encountered during training, and (3) their robustness to trojans implanted during pretraining. We find that LAT in the optimal layer usually performs better than AT under all three evaluations.
@@
    We show in vision, language understanding, and language generation tasks that LAT can improve robustness against trojans and novel classes of attacks. We find that LAT in the optimal layer usually Pareto dominates AT with respect to both clean and robust performance.
@@
    We find cautionary instances in which robustness techniques sometimes harm robustness. We show an instance in which $L_{p}$-norm AT in vision models makes a network more susceptible to novel attacks. Also similar to recent findings from Hubinger et al. (2024), we show instances in which, under certain configurations, AT and LAT (without using examples containing a trojan trigger) can make a trojaned LLM more susceptible to its trojan.
@@ 2 Related Work
Unforeseen failure modes and empirical shortcomings of adversarial training: Some problems with modern deep learning systems can be discovered using test sets, red-teaming, or adversarial attacks. In these cases, practitioners can apply AT and other techniques to address these failures before deployment (Madry et al., 2017; Achiam et al., 2023; Ganguli et al., 2022; Anil et al., 2023; Touvron et al., 2023). However, some failures that are hard to find during development can still appear post-deployment (Hendrycks et al., 2021a; Goel et al., 2024).
  • Trojans (also known as backdoors) can be triggered by arbitrary features that are hard for developers to find (Chen et al., 2017; Wu et al., 2022; Carlini et al., 2023).
  • Jailbreaks can elicit harmful outputs from language models subverting safety constraints. Jailbreaking prompts can take a variety of forms including unintelligible text (Zou et al., 2023b), persona modulation (Shah et al., 2023), ASCII art (Jiang et al., 2024), low-resource languages (Yong et al., 2023), encoded prompts (Wei et al., 2023), and other strategies (Shen et al., 2023; Liu et al., 2023; Rao et al., 2023).
  • Other novel attacks aside from jailbreaks, have also elicited unwanted behaviors from AI systems (Brown et al., 2018; Shayegani et al., 2023; Geiping et al., 2024).
  • Black swans refer to anomalous failure modes which can avoid detection due to their rarity (Kolt, 2023; Hendrycks et al., 2021b).
Recently, the deployment of modern AI systems has set off ongoing games of ‘cat-and-mouse’ in which developers continually update their models in response to newly discovered vulnerabilities (e.g., (at DEF CON, 2023; OpenAI, 2023)).
Limitations of (adversarial) fine-tuning’s ability to generalize: AT generally requires examples of a failure in order to fix it. Hubinger et al. (2024) and Jain et al. (2023a) have both shown cases in which AT can fail to fix specific problems with LLMs that occur off the attack distribution. Ziegler et al. (2022) also found that adversarially trained language classifiers remained somewhat vulnerable to the same attack-generation method as was used during training. These shortcomings can be explained in part by memorization or “shortcut learning” of spurious features instead of the desired higher-level concepts (Geirhos et al., 2020; Du et al., 2023).
@@
Limitations of fine-tuning for making mechanistic changes in language models: Standard fine-tuning does not directly shape a model’s inner knowledge or representations – it only directly supervises or reinforces its behavior. However, rarely-used latent capabilities inside models can cause harm if they resurface (e.g., via a jailbreak). It has been found that undesirable dormant capabilities can be elicited from LLMs by pruning (Wei et al., 2024) and few-shot fine-tuning (Yang et al., 2023; Qi et al., 2023; Lermen et al., 2023; Zhan et al., 2023). For LLMs, these results are relatively unsurprising in light of recent work showing that they resist forgetting (Ramasesh et al., 2021; Cossu et al., 2022; Li et al., 2022; Scialom et al., 2022; Luo et al., 2023; Kotha et al., 2023; Shi et al., 2023) and that fine-tuning struggles to make major changes to a model’s inner knowledge and capabilities (Lubana et al., 2023; Juneja et al., 2022; Jain et al., 2023b; Lee et al., 2024; Prakash et al., 2024). Jain et al. (2023b) likened fine-tuning in LLMs to merely modifying a “wrapper” around a stable, general-purpose set of latent capabilities.
Latent space attacks in language models: In language models, it is not possible to directly use gradient-based methods to generate adversarial attacks because tokenization is not differentiable. However, several works have attacked word embeddings (which can be viewed as the first latent state in the network) and trained on these perturbations to improve robustness or generalization (Jiang et al., 2019; Zhu et al., 2019; Liu et al., 2020; He et al., 2020; Kuang and Bharti, ; Li and Qiu, 2021; Sae-Lim and Phoomvuthisarn, 2022; Pan et al., 2022; Schwinn et al., 2023; Geisler et al., 2024; Schwinn et al., 2024). Here, we use embedding space adversarial training as a baseline. Deeper in the latent space, Fort (2023) conducted simple experiments to suggest that language models can be very sensitive to latent perturbations. Others have shown that LLMs can have their high-level behaviors altered by perturbations to their latent states found through probing or causal interventions (Turner et al., 2023; Li et al., 2023; Zou et al., 2023a; Wang and Shu, 2023; Rimsky et al., 2023; Jorgensen et al., 2023; Lu and Rimsky, 2024; von Rütte et al., 2024). However, to the best of our knowledge, these types of perturbations have not yet been used for LAT. The most similar work to ours is Kitada and Iyatomi (2023), who perform LAT on attention representations to improve generalization in small BERT-scale models. In contrast, however, we use models up to 7B parameters in size, use more general attacks by targeting the residual stream of transformers, test both classifiers and generators, and evaluate methods under trojans and novel classes of adversarial attacks.
@@ 3 Method
Threat model: The threat that we consider is not an attacker that has access to the latent space. Although we train under latent adversarial perturbations, our ultimate goal is not to make the trained model resistant to these. Instead, our goal is to make models robust to distribution shifts between development and deployment that are not precisely known beforehand (e.g., trojans, jailbreaks, novel attacks, and black swans).
@@
Both the inner and outer optimization problems are typically solved using gradient-based optimization on $\delta_{i}^{x}$ and $\theta$, respectively.
@@
Note that this setup involves ‘untargeted’ attacks in which the adversary aims to maximize the target model’s loss. ‘Targeted’ attacks in which the adversary attempts to elicit a particular harmful output are also possible, but we leave this to future work.
In our experiments, for all attacks, we also clip the perturbed activations by the min and max of the unperturbed activations across all neurons in the current batch to reduce the risk of attacks moving activations to an irrelevant part of the latent space.
Latent space distance metric: Typically, AT constrains the perturbation according to a constraint defined by a simple distance metric such as an $L_{p}$-norm. This is reasonable for AT because all input components (e.g., pixels) have comparable activation distributions. However, this is not guaranteed for latent representations – each neuron may have a different distribution of activations. As a result, we experiment with using a normalized distance metric to constrain latent adversarial perturbations. However, we find no clear differences in results between using standard and normalized distance metrics. In Section 4, we pool results using standard and normalized distance metrics together, but in Appendix A, we present each side-by-side.
@@ 4 Experiments
We experiment with three different tasks: image classification on ImageNet (Russakovsky et al., 2014), text classification on the Redwood injurious text dataset (Ziegler et al., 2022), and text generation on the Anthropic Helpful-Harmless-RLHF (Bai et al., 2022) and PKU BeaverTails (Ji et al., 2024) data distributions. In each experiment, we compare 3 methods: AT, LAT, and training under random latent perturbations (RLP). We use this as a non-adversarial contrast to LAT. We select the latent layer to perturb by performing preliminary experiments with LAT across layers and selecting one in which it performed well overall. We produce all attacks using projected gradient descent (Madry et al., 2017).
For each experiment, we first fine-tune the model on the target task using poisoned data to implant trojans. Second, we fine-tune further on clean data while applying RLP, AT, and LAT. We report results across this second fine-tuning stage. For each task, we evaluate methods based on (1) performance on clean data, (2) robustness to novel classes of input-space adversarial attacks, and (3) robustness to the trojans we implanted. We illustrate this in Figure 1d.
Because in different applications, practitioners may prefer different tradeoffs between clean and robust performance, we focus on the Pareto frontier between clean performance and each type of robustness. In each experiment, we perform multiple runs of RLP, AT, and LAT with varying perturbation bounds ($\epsilon$) and evaluate the model at multiple training checkpoints for each. This allows us to construct a scatterplot of different tradeoffs between clean and robust performance. Overall, we find that LAT in the optimal layer usually Pareto dominates RLP and AT with respect to both clean and robust performance.
@@ 4.1 Image Classification
We used a ResNet-50 from He et al. (2016) and fine-tuned it on a version of the ImageNet (Russakovsky et al., 2014) training set that was poisoned as in Casper et al. (2023) to implant 8 trojans: four with a $64\times 64$ pixel patch trigger, and four with a natural feature trigger which took the form of a pre-existing feature in certain images. In both cases, the trojan response was a randomly selected label different from the ground truth. Based on preliminary experiments, we selected the activations after the residual blocks (pre-avgpooling) to perturb for LAT. We then fine-tuned the model on clean ImageNet training data for one epoch using RLP, AT, and LAT. Finally, we evaluated the resulting models on (1) the clean ImageNet validation set, (2) 18 assorted unseen attack methods from Kaufmann et al. (2019), [^3] We used 18 of the 20 attacks from Kaufmann et al. (2019) excluding FGSM and PGD. and (3) images with the 8 trojan triggers.
We plot (1) vs (2) and (1) vs (3) in Figure 2. All Pareto frontiers are entirely filled with model checkpoints from LAT. We also find an example of how $L_{p}$-norm AT can be harmful to robustness against novel attacks (see Figure 2a and Figure 5a). AT using $L_{p}$-norm attacks caused the model to be more susceptible to other attacks from Kaufmann et al. (2019). [^4] Kaufmann et al. (2019) found that $L_{p}$ norm AT could be used to improve robustness to their attacks. Our results do not conflict with this. We find in Figure 2 that some checkpoints and some runs of AT did improve robustness, but we nonetheless found that most checkpoints from most AT runs were less robust than before AT.

### v4

@@ Abstract
Despite extensive diagnostics and debugging by developers, AI systems sometimes exhibit harmful unintended behaviors. Finding and fixing these is challenging because the attack surface is so large – it is not tractable to exhaustively search for inputs that may elicit harmful behaviors. Red-teaming and adversarial training (AT) are commonly used to improve robustness, however, they empirically struggle to fix failure modes that differ from the attacks used during training. In this work, we utilize latent adversarial training (LAT) to defend against vulnerabilities without leveraging knowledge of what they are or using inputs that elicit them. LAT makes use of the compressed, abstract, and structured latent representations of concepts that the network actually uses for prediction. Here, we use it to defend against failure modes without examples that elicit them. Specifically, we use LAT to remove trojans and defend against held-out classes of adversarial attacks. We show in image classification, text classification, and text generation tasks that LAT usually improves both robustness to novel attacks and performance on clean data relative to AT. This suggests that LAT can be a promising tool for defending against failure modes that are not explicitly identified by developers. [^1] Code is available at https://github.com/thestephencasper/latent_adversarial_training.
@@ 1 Introduction
Ensuring that AI systems will be trustworthy, even in the face of anomalous and adversarial inputs, has been a major focus of research in the past decade (Szegedy et al., 2013; Goodfellow et al., 2014; Zhao et al., 2022; Anwar et al., 2024; Yohsua et al., 2024), and it has been incorporated into risk management frameworks for AI governance (NIST, 2023; DSIT, 2023; EU, 2021; NISSTC, 2023). Developers commonly use test sets, red-teaming, and attack methods to identify vulnerabilities followed by adversarial training (AT) to fix them. This is valuable, but sometimes fails to address problems. There are often systematic differences between the failure modes that developers search for (e.g., $L_{p}$-norm attacks) and ones that models can exhibit post-deployment (e.g. (Hendrycks et al., 2021a, c)). Many real-world vulnerabilities can evade detection such as trojans (Hubinger et al., 2024; Carlini et al., 2022), jailbreaks (Liu et al., 2023; Wei et al., 2023; Zou et al., 2023b; Shah et al., 2023) novel attacks (Brown et al., 2018; Shayegani et al., 2023; Geiping et al., 2024), or black swans (Kolt, 2023; Hendrycks et al., 2021b). In Figure 1a, we illustrate the gap between failure modes developers identify, and the “unforeseen” ones they do not.
@@
Our motivation for LAT is based on the vastness of a model’s input space compared to the manifold of task-relevant features within it. Across the latents, a model gradually develops more compressed, abstract, and structured representations of the concepts it uses to process information. [^2] Since the mappings from layer to layer are not typically surjective, not all latent states are inducible by real inputs. This relaxation of the problem is what gives LAT the potential to compensate for the gap between failure modes that are identified by developers and ones that are not (see Figure 1a). This makes it possible for latent space attacks to activate neural circuitry that elicits failures (Figure 1b) without requiring inputs that trigger them (Fort, 2023). Thus, we hypothesize that even if it is difficult to find a problem with a model using input space attacks, it may be comparatively easier with latent space ones. We make three contributions:
Figure 1: (a) We study latent adversarial training (LAT) as a method to reduce risks from failure modes that are not identified by developers pre-deployment. (b) Our motivation for LAT is based on how models develop more compressed, abstract, and structured representations across their latents. We hypothesize that many failures that are difficult to elicit from the input space may be easier to elicit from the latent space. (c) In experiments, we compare AT with LAT. The key difference is that AT applies adversarial perturbations to the input, and LAT applies adversarial perturbations to a hidden layer. (d) In each of our experiments, we compare methods based on (1) their performance on clean evaluation data, (2) their robustness to novel classes of adversarial examples not encountered during training, and (3) their robustness to trojans implanted during pretraining. We find that LAT in the optimal layer usually performs better than AT under all three evaluations.
@@
    We show in vision, language understanding, and language generation tasks that LAT can improve robustness against failure modes without any examples that elicit them. Specifically, we use it to defend against trojans and novel classes of attacks. We find that LAT in the optimal layer usually Pareto dominates AT with respect to both clean and robust performance.
@@
    We demonstrate cautionary instances in which robustness techniques sometimes harm robustness to novel failure modes. We show an instance in which $L_{p}$-norm AT in vision models makes a network more susceptible to novel attacks. Also similar to recent findings from Hubinger et al. (2024), we show instances in which, under certain configurations, AT and LAT (without using examples containing a trojan trigger) can make a trojaned LLM more susceptible to its trojan.
@@ 2 Related Work
Unforeseen failure modes and empirical shortcomings of adversarial training: Some problems with modern deep learning systems can be discovered using test sets, red-teaming, or adversarial attacks. In these cases, practitioners can apply AT and other techniques to address these failures (Madry et al., 2017; Achiam et al., 2023; Ganguli et al., 2022; Anil et al., 2023; Touvron et al., 2023). However, some failures that are hard to find during development can still appear post-deployment (Hendrycks et al., 2021a; Goel et al., 2024; Carlini et al., 2024).
  • Trojans (also known as backdoors) can be triggered by arbitrary features (Chen et al., 2017; Wu et al., 2022; Carlini et al., 2023).
  • Jailbreaks can elicit harmful outputs from language models subverting safety constraints. Jailbreaking prompts can take a variety of forms including persuasive text (Shen et al., 2023; Liu et al., 2023), persona modulation (Shah et al., 2023), low-resource languages (Yong et al., 2023), long-context attacks (Anil et al., 2024), encoded prompts (Wei et al., 2023), unintelligible text (Zou et al., 2023b), ASCII art (Jiang et al., 2024), images (Bailey et al., 2023), and other strategies (Shen et al., 2023; Rao et al., 2023; Andriushchenko et al., 2024).
  • Other novel attacks aside from jailbreaks, can elicit unwanted behaviors from AI systems (Brown et al., 2018; Laidlaw et al., 2020; Dai et al., 2022; Shayegani et al., 2023; Geiping et al., 2024; Laidlaw et al., 2020; Chang et al., 2024).
  • Black swans refer to anomalies which avoid detection due to their rarity (Kolt, 2023; Hendrycks et al., 2021b).
Recently, the deployment of modern AI systems has set off ongoing games of ‘cat-and-mouse’ in which developers continually update their models in response to newly discovered exploits.
Limitations of (adversarial) fine-tuning’s ability to generalize: AT generally requires examples of a failure in order to fix it. Hubinger et al. (2024) and Jain et al. (2023a) have both shown cases in which AT can fail to fix specific problems with LLMs that occur off the attack distribution. Ziegler et al. (2022) also found that adversarially trained language classifiers remained somewhat vulnerable to the same attack-generation method used during training. These shortcomings can be explained in part by memorization or “shortcut learning” of spurious features instead of the desired concepts (Geirhos et al., 2020; Du et al., 2023).
@@
Limitations of fine-tuning for making mechanistic changes in language models: Standard fine-tuning does not directly shape a model’s inner knowledge or representations – it only directly supervises or reinforces its behavior. However, rarely-used latent capabilities can cause harm if they resurface (e.g., via a jailbreak). Undesirable dormant capabilities can be elicited from LLMs by pruning (Wei et al., 2024) and few-shot fine-tuning (Yang et al., 2023; Qi et al., 2023; Lermen et al., 2023; Zhan et al., 2023; Wei et al., 2024). For LLMs, these results are relatively unsurprising in light of recent work showing that they resist forgetting (Ramasesh et al., 2021; Cossu et al., 2022; Li et al., 2022; Scialom et al., 2022; Luo et al., 2023; Kotha et al., 2023; Shi et al., 2023) and that fine-tuning struggles to make major changes to latent knowledge and capabilities (Lubana et al., 2023; Juneja et al., 2022; Jain et al., 2023b; Lee et al., 2024; Prakash et al., 2024; Qi et al., 2024). Jain et al. (2023b) likened fine-tuning in LLMs to merely modifying a “wrapper” around a stable, general-purpose set of latent capabilities.
Latent space attacks in language models: In language models, it is not possible to directly use gradient-based methods to generate adversarial attacks because tokenization is not differentiable. However, several works have attacked word embeddings (which can be viewed as the first latent state in the network) and trained on these perturbations to improve robustness or generalization (Jiang et al., 2019; Zhu et al., 2019; Liu et al., 2020; He et al., 2020; Kuang and Bharti, ; Li and Qiu, 2021; Sae-Lim and Phoomvuthisarn, 2022; Pan et al., 2022; Schwinn et al., 2023; Geisler et al., 2024; Schwinn et al., 2024; Xhonneux et al., 2024). Here, we use embedding space adversarial training as a baseline. Deeper in the latent space, Fort (2023) demonstrated that language models can be very sensitive to latent perturbations. Others have shown that LLMs can have their high-level behaviors altered by perturbations to their latent staftes found through probing or causal interventions (Turner et al., 2023; Li et al., 2023; Zou et al., 2023a; Wang and Shu, 2023; Rimsky et al., 2023; Jorgensen et al., 2023; Lu and Rimsky, 2024; von Rütte et al., 2024). However, to the best of our knowledge, these types of perturbations have not yet been used for LAT. The most similar work to ours is Kitada and Iyatomi (2023), who perform LAT on attention representations to improve generalization in small BERT-scale models. In contrast, however, we use models up to 7B parameters in size, use more general attacks by targeting the residual stream of transformers, test both classifiers and generators, and evaluate methods under trojans and novel attacks.
@@ 3 Method
Threat model: The threat that we consider is not an attacker that has access to the latent space. Although we train under latent perturbations, our ultimate goal is not to make the trained model resistant to these. Instead, our goal is to make models robust to distribution shifts between development and deployment that are not precisely known beforehand (e.g., trojans, jailbreaks, novel attacks, and black swans).
@@
Both the inner and outer problems are typically solved with gradient-based optimization on $\delta_{i}^{x}$ and $\theta$, respectively.
@@
Note that this setup involves ‘untargeted’ attacks in which the adversary maximizes the target model’s loss. ‘Targeted’ attacks in which the adversary elicits a specific target output are possible, but we leave this to future work.
For all attacks, we clip the perturbed activations by the min and max of the unperturbed activations across all neurons in the current batch to reduce the risk of attacks moving activations to an irrelevant part of the latent space.
Latent space distance metric: Typically, AT constrains the perturbation according to a constraint defined by a simple distance metric such as an $L_{p}$-norm. This is reasonable for AT because all input components (e.g., pixels) have comparable activation distributions. However, this is not guaranteed for latent representations – each neuron may have a different distribution of activations. As a result, we experiment with using a normalized distance metric to constrain latent perturbations. However, we find no clear differences in results between using standard and normalized distance metrics. In Section 4, we pool results using standard and normalized distance metrics together, but in the Appendix, we present each side-by-side.
@@ 4 Experiments
Three tasks: image classification, text classification, and text generation. We experiment with three different tasks: image classification on ImageNet (Russakovsky et al., 2014), text classification on the Redwood injurious text dataset (Ziegler et al., 2022), and text generation on the Anthropic Helpful-Harmless-RLHF (Bai et al., 2022) and PKU BeaverTails (Ji et al., 2024) data distributions.
Three methods: AT, LAT, and RLP. In each experiment, we compare 3 methods: AT, LAT, and training under random latent perturbations (RLP). We use these random latent perturbations with the same norm as LAT perturbations as a non-adversarial contrast to LAT. We select the latent layer to perturb by sweeping across layers for high clean and robust performance. We converged to the heuristics of perturbing the first post-convolutional layer in CNNs and a relatively early layer in transformers. We produce all attacks using projected gradient descent (Madry et al., 2017).
Three measures: clean performance, adversarial robustness, and trojan forgetting. For each experiment, we first fine-tune the model using poisoned data to implant trojans. Second, we fine-tune further on clean data while applying RLP, AT, and LAT. We report results across this second fine-tuning stage. For each task, we evaluate methods based on (1) performance on clean data, (2) robustness to novel classes of input-space adversarial attacks, and (3) robustness to trojans we implanted. We illustrate this in Figure 1d.
Our goal: Expanding the Pareto frontier. Because in different applications, practitioners may prefer different tradeoffs between clean and robust performance, we focus on the Pareto frontier between clean performance and each type of robustness. In each experiment, we perform multiple runs of RLP, AT, and LAT with varying perturbation bounds ($\epsilon$) and evaluate at multiple training checkpoints for each. This allows us to construct a scatterplot of different tradeoffs between clean and robust performance. Overall, we find that LAT in the optimal layer usually Pareto dominates RLP and AT with respect to both clean and robust performance.
@@ 4.1 Image Classification
We used a ResNet-50 from He et al. (2016) and fine-tuned it on a version of the ImageNet (Russakovsky et al., 2014) training set that was poisoned as in Casper et al. (2023) to implant 8 trojans: four with a $64\times 64$ pixel patch trigger, and four with a natural feature trigger which took the form of a pre-existing feature in certain images. In both cases, the trojan response was a randomly selected label different from the ground truth. We then fine-tuned the model on clean ImageNet training data for one epoch using RLP, AT, and LAT. When performing LAT, we attacked the first layer activations after the residual blocks (pre-avgpooling). Finally, we evaluated the resulting models on (1) the clean ImageNet test set, (2) the ImageNet test set images, each attacked using one of the 18 held-out attack methods from Kaufmann et al. (2019), [^3] We used 18 of the 20 attacks from Kaufmann et al. (2019) excluding FGSM and PGD to avoid contaminating the test set with $L_{p}$-norm input-space attacks. and (3) the ImageNet test set images, each attacked with one of the 8 trojan triggers.
We plot (1) vs (2) and (1) vs (3) in Figure 2. All Pareto frontiers are entirely filled with model checkpoints from LAT. Compared to AT, LAT methods result in 68%, 68%, and 1.5% greater improvements to the area under the Pareto curve for attack robustness (a) patch trojan removal (b.1) and natural feature trojan removal (b.2) respectively. We also find an example of how $L_{p}$-norm AT can be harmful to robustness against novel attacks (see Figure 2a and Figure 5a). AT using $L_{p}$-norm attacks caused the model to be more susceptible to other attacks from Kaufmann et al. (2019). [^4] Kaufmann et al. (2019) found that $L_{p}$ norm AT could be used to improve robustness to their attacks. Our results do not conflict with this. We find in Figure 2 that some checkpoints and some runs of AT did improve robustness, but we nonetheless found that most checkpoints from most AT runs were less robust than before AT.